Calculate using Scaffolding method and explain each steps 265÷7

589÷4

735÷3

4581÷7


1
Expert's answer
2021-07-26T17:21:43-0400

265÷7

"30\cdot 7=210" , remainder is 55

"7\cdot 7=49" , remainder is 6

Answer: "30+7=37" , remainder is 6


589÷4

"100\cdot 4=400" , remainder is 189

"40\cdot 4=160" , remainder is 29

"7\cdot 4=28" , remainder is 1

Answer: "100+40+7=147" , remainder is 1


735÷3

"200\cdot 3=600" , remainder is 135

"40\cdot 3=120" , remainder is 15

"5\cdot 3=15" , remainder is 189

Answer: "200+40+5=245"


4581÷7

"600\cdot 7=4200" , remainder is 381

"50\cdot 7=350" , remainder is 31

"4\cdot 7=28" , remainder is 3

Answer: "600+50+4=654" , remainder is 3